Consumers consumption (C) = 160$

Investment (I) = 40$

Government expenditure (G) = 30$

Export (X) = 20$

Import (M) = 10$

Indirect taxes = 25$

Subsidies = 118$

GDP Market Price = 160$ + 40$ + 30$ + (20$ - 10$) = 230$

= 230$ + 10$ = 240$

GDP Factor Cost = 240$ - 25$ + 118$ = 333$

it my answer correct ? If wrong please help me to correct sorry to bother you thank you so much.

1
Expert's answer
2021-01-25T14:22:17-0500

Y= C+I+G+(X-M)

Where Y is the GDP

C personal consumption expenditure

I gross private domestic investment

G government expenditure

(X-M) net export

X is exports and M imports

Y=160+40+30+(20-10)

Y=230+10

Y=240&at market price

GDP at factor cost=GDP at market price - depreciation- indirect taxes + subsidies

Y=240-0-25+118

Y=333$ at factor cost

your answer is correct